### v1

@@ 1 Fabricating functional suspended devices on silicon-on-insulator
A critical step to achieve post-release lithography was the development of a method to support the released devices, embedding them in partially solidified photoresist (Fig. 1). For this step, the photoresist was diluted with edge bead removal solvent to reduce its viscosity and allow gentle spin coating at a low speed of 1000 RPM. With an empirically determined dilution ratio of equal parts photoresist and solvent, we found that this process did not harm the released structures, provided sufficient support during the subsequent fabrication procedures, and allowed easy removal during the final liftoff of the excess magnetostrictive material. Attempts using undiluted photoresist, with its relatively high viscosity and requiring typical spin coating speeds, broke the structures.

### v2

@@ 1 Fabricating functional suspended devices on silicon-on-insulator
A critical step to achieve post-release lithography was the development of a method to support the released devices, embedding them in partially solidified photoresist (Fig. 1). For this step, the photoresist was diluted with edge bead removal solvent with an equal volume of edge bead removal solvent, reducing its viscosity by a factor of five. Combined with the factor of four reduction in spin speed, and including the change in resist thickness due to the modified viscosity, this results in a factor of twenty lower shear load during spinning compared to undiluted resist. The diluted resist acts as an elastic support for the suspended device and can be modeled as an elastic Winkler Foundation [hetenyi1946beams]. Within this model, the resist exerts a distributed restoring force, with characteristic length µm, where , , is the Young’s modulus of the suspended structure with width and thickness , and is the Young’s modulus of the resist with thickness . Because the length of the suspended structure is much larger than the characteristic length, the foundation exerts strong restraint on the beam and constrains out-of-plane motion, suppressing global bending and thereby protecting the delicate released structures during subsequent fabrication steps. We found that this post-release lithography process did not harm the released structures, provided sufficient support and allowed easy removal during the final liftoff of the excess magnetostrictive material. Attempts using undiluted photoresist, with its relatively high viscosity and requiring typical spin coating speeds, broke the structures.